OpenStudy (anonymous):

A rocket is launched with acceleration of "a ft/sec^2". Neglecting the air resistance, its height after t second is given by f(a, t) = 1/2 (a-32)t^2 feet. Fuel used by the rocket is proportional to a^2 t and the limited fuel capacity of the rocket follows the equation a^2 t = 10000. Find the accelration so that the rocket will reach maximum height when the fuel runs out. What is this height?

OpenStudy (anonymous):

Well, first we can express the acceleration as a function of time: \[a=(\frac{10^4}{t})^{\frac{1}{2}}=\frac{100}{\sqrt{t}}\]If we plug this into our equation for height, we get a function of t alone: \[f(t)=\frac{1}{2}(\frac{100}{\sqrt{t}}-32)t^2=50t^{\frac{3}{2}}-16t^2\]Now we are asked to maximize this height, so we find the max value of this function with respect to t. Once we have this value of t, the acceleration follows immediately from the relation a^2t=10000

OpenStudy (anonymous):

good from here?

OpenStudy (anonymous):

not so much! i understood till what you did. What do i need to do after that?

OpenStudy (anonymous):

You need to use your calculus techniques to find the max value of this height function

OpenStudy (anonymous):

yea. Got it. Thanks

OpenStudy (anonymous):

That will give you the height. Whatever t is equal to at this height is used to find the acceleration of the rocket.

OpenStudy (anonymous):

so do i do t = 10000/a^2

OpenStudy (anonymous):

I get t=5.5s (rounded), giving height=161 ft (rounded) and acceleration=42.7 ft/s^2 (rounded)

OpenStudy (anonymous):

Once you get t, you use:\[a=\sqrt{ \frac{10000}{t}}\] to get a.

OpenStudy (anonymous):

oooops...I used the wrong expression for a

OpenStudy (anonymous):

looks like \[a^{2t}=10000\]I was using\[at^2=10000\]

OpenStudy (anonymous):

no no.. its a^2 * t = 10000

OpenStudy (anonymous):

so a^2 = 10000/t and then a = sqrt( 10000/t)

OpenStudy (anonymous):

eeeer....ok. nm...I did it correct then. I did use \[a^2t=10000\]

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

any questions?

OpenStudy (anonymous):

off to play call of duty...good luck

OpenStudy (anonymous):

I am struggling to find t

OpenStudy (anonymous):

\[f'(t)=\frac{150}{2}t^\frac{1}{2}-32t=0\]\[\sqrt{t}= \frac{150}{64}\]So,\[t \approx5.5\]

OpenStudy (anonymous):

awesome and i did get a as 42.7 and max height at 161 rounded. Now as an additional information what do i do to find the value of a that maximizes the total height of the rocket. And then for this, what is the maximum height. How do i proceed with that?

OpenStudy (anonymous):

fuel is a limitation. use lagrange multiplier

OpenStudy (anonymous):

and can you show me how to do that? please. :(

OpenStudy (anonymous):

just wondering. you know what lagrange multiplier is right?

OpenStudy (anonymous):

yea its used to fine max/min values.

OpenStudy (anonymous):

k then ill show u the first few steps

OpenStudy (anonymous):

hold on. will take a while to draw out

OpenStudy (anonymous):